OpenStudy (anonymous):

graph -2x+4y=4 for the domain D:{-8,-4,0,4,8}

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

are you asking if it is righ or what

OpenStudy (anonymous):

no thats the question

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

do you have to type in the anwser or is there muliple anwsers

OpenStudy (anonymous):

u have to type

OpenStudy (anonymous):

okay i thing that this is it y = 3x -13

OpenStudy (anonymous):

how did u figure that out

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

1. (a) Is the graph of a circle a graph of a function? Yes or no (b) Find the center and the radius of the circle represented by the equation x^2 + y^2 + 10x – 4y – 7 = 0 x^2+10x+25 + y^2 -4y + 4 = 7+25+4 (x+5)^2 + (y-2)^2 = 36 ----- Center: (-5,2) ; radius = 6 ----------- ----------- -------- 2. Let g(x)= 9-x/6+x (a) Find g(–1) g(-1) = (9 + 1)/(6-1) = 2 ----------- ----------- -------- (b) Find g(t + 1) and simplify. g(t+1) = (9-(t+1))/(6+(t+1)) g(t+1) = (8-t)/(7+t) ----------- ----------- ------- (c) Find the domain of the function Let g(x)= 9-x/6+x Domain: All Real Numbers except x=-6

OpenStudy (anonymous):

3.Consider the points (4, –1) and (7, 8). (a) Find the midpoint of the line segment with the given endpoints. x = (4+7)/2 = 11/2 y = (-1+8)/2 = 7/2 ----------- ----------- - (b) Find the distance between the points. Give an exact answer (simplified as much as possible), and also an approximation to three decimal places. d = sqrt[(8--1)^2+(7-4)^2] = sqrt(117) ----------- ----------- ----------- ------ (c) Find the slope of the line passing through the two given points slope = (8--1)/(7-4) = 9/3 = 3 ----------- --------- (d) Find the equation of the line passing through the two given points. 8 = 3*7 + b b = -13 Anwser is y = 3x -13

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

ur welcome